# CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents

### v4

@@ Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets face limitations: they rely on human annotators who assess code primarily through semantic understanding rather than functional verification, leading to potential inaccuracies and scalability issues. Additionally, current evaluation metrics often overlook the multi-choice nature of code search. This paper introduces CoSQA+, pairing high-quality queries from CoSQA with multiple suitable codes. We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system. Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 96.4%. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
[…]
tcb@breakable
CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents
Jing Gongthanks: Equal Contribution, Yanghui Wu11footnotemark: 1, Linxi Liang11footnotemark: 1, Yanlin Wangthanks: Corresponding Author, Jiachi Chen, Mingwei Liu, Zibin Zheng Sun Yat-sen University {gonj39,wuyh223,lianglx26}@mail2.sysu.edu.cn, yanlin-wang@outlook.com, {chenjch86,liumw26,zhzibin}@mail.sysu.edu.cn,
Introduction
@@
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development  (Li et al., 2013; Sim et al., 2011; Stolee et al., 2014). Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems  (Yan et al., 2020). Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR)  (Husain et al., 2019; Huang et al., 2021; Li et al., 2024a; Lu et al., 2021; Yao et al., 2018; Feng et al., 2020). According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average. For each query, they typically reference 2.64 code examples to find the necessary information. These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
Existing code search datasets like CoSQA  (Huang et al., 2021) and CodeSearchNet  (Husain et al., 2019) rely on expert human annotators to evaluate query-code pairs. While this approach captures human judgment, it exists several limitations. First, human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations. Second, manual annotation requires significant time and resources, limiting dataset scalability. Recent works by  Li et al. (2024a) leverage LLMs for annotation to address scalability issues. However, LLMs face challenges in reliably determining whether code functionally fully satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic. Furthermore, most code search benchmarks employ metrics like MRR or Pass@k (Lu et al., 2021), which primarily evaluate single-result retrieval performance. This approach overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations. While alternative metrics like MAP and Recall@k Husain et al. (2019) exist for evaluating multiple code search, they underemphasize the importance of top-ranked results. This is problematic since, according to our survey, users heavily rely on first-page search results.
@@
In this work, we begin with a survey of 102 experienced Python programmers to understand multi-choice code search practices. The survey results reveal that developers frequently require and utilize multiple code examples for a query. Motivated by these insights, we present a new benchmark CoSQA+, with 373,339 pairs of web queries filtered from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search. To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a first screener, test program generator, test executor, bug fixer, and final arbiter. The label indicates whether each code snippet exactly matches its corresponding query requirements. Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks  (Wang et al., 2013). We use it to measure 9 code search methods, 8 of which are embedding-based.
@@
To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 200 randomly selected query-code pairs from CoSQA+. Three Python experts first design test programs for these pairs to establish ground truth annotations. We then compare the accuracies among test-driven agents, python expert annotators without tests, and single LLM. Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 96.4%. Notably, 83.67% of the test programs automatically generated by test-driven agents are executable. Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings. Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both NDCG@10 and MRR metrics.
@@
  • To the best of our knowledge, we develop the first Multiagent annotator with over 96.4% accuracy for code search dataset.
@@
Related Work
Code Search Datasets
Besides our work, there are other datasets and benchmarks in the realm of code search.  Husain et al. (2019) present CodeSearchNet which contains a wide range of query-code pairs from GitHub. CoSQA  (Huang et al., 2021) and CS1QA  (Lee et al., 2022) contain the search logs collected from Microsoft Bing search engine but only cover Python and are limited in size. StaQC  (Yao et al., 2018) is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality.  Li et al. (2024b) introduces ProCQA, a large-scale community-based programming question answering dataset mined from StackOverflow with strict filtering strategies for quality and fairness control. This benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague. Recently,  Li et al. (2024a) present Query4Code, which also leverages LLMs for dataset annotation, but their approach uses a single model for similarity computation and adopts a less strict matching criterion.
Test-Driven Agents
@@
Traditional test generation approaches relied heavily on human expertise, but recent work has demonstrated the potential of automated test agents. LLM-based agents have shown promising results in various testing scenarios. For instance, in the domain of deep learning libraries,  Deng et al. (2024) develops agents capable of generating diverse and meaningful test inputs. Several studies have explored LLM-based agents for specialized tasks. For test oracle generation,  Nashid et al. (2023) utilizes few-shot prompting to create context-aware assertions, achieving a high accuracy. In program repair, studies by  Xia and Zhang (2023) and  Jiang et al. (2023) have demonstrated the effectiveness of integrating LLMs with iterative validation workflows. However, existing test-driven agents primarily focus on validating specific code functionality without addressing the semantic challenges in code search systems, where our test-driven annotation approach bridges this gap through systematic validation of query-code relevance.
Figure 1: The construction process of CoSQA+. The final CoSQA+ dataset includes queries from original CoSQA, and codes filtered from CodeSearchNet.
CoSQA+
@@
Figure 1 illustrates the benchmark construction process for CoSQA+. The process starts with query and code collection, in which we leverage the queries from CoSQA and collect python code snippets filtered from CodeSearchNet. In candidate pairs construction, collected code snippets and queries are embedded as vectors to calculate similarity. For each query, the top 20 query-code pairs with the highest similarity are selected as candidates for potential matches. In model annotation, each query-code pair is annotated by test-driven agents with “1” or “0”, indicating whether the code snippet exactly matches the query. These labeled pairs are then added to CoSQA+.
Query and Code Collection
@@
This section details the reasons why we select CoSQA as the query source and the CodeSearchNet Corpus  (Husain et al., 2019) serves as the code sources.
@@
Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes  (Huang et al., 2021). We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns. For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets  (Sun et al., 2022). (2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent. This careful curation helps maintain high data quality and reduces noise in the training process  (Li et al., 2020). By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
@@
To the best of our knowledge, CodeSearchNet is the only origin dataset specifically designed for Python functions in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages  (Husain et al., 2019). According to our survey, code searchers prefer function codes with clear input and output. These codes are high quality with higher readability, testability and maintainability  (Martin, 2009). To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach. This involves parsing the code into an Abstract Syntax Tree (AST), identifying function nodes with at least one input parameter and a valid return statement. Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
Candidate Query-Code Pairs Construction

### v5

# CoSQA+: Enhancing Code Search Evaluation with a Multi-Choice Benchmark and Test-Driven Agents

@@ Abstract
Semantic code search, retrieving code that matches a given natural language query, is an important task to improve productivity in software engineering. Existing code search datasets face limitations: they rely on human annotators who assess code primarily through semantic understanding rather than functional verification, leading to potential inaccuracies and scalability issues. Additionally, current evaluation metrics often overlook the multi-choice nature of code search. This paper introduces CoSQA+, pairing high-quality queries from CoSQA with multiple suitable codes. We develop an automated pipeline featuring multiple model-based candidate selections and the novel test-driven agent annotation system. Among a single Large Language Model (LLM) annotator and Python expert annotators (without test-based verification), agents leverage test-based verification and achieve the highest accuracy of 92.0%. Through extensive experiments, CoSQA+ has demonstrated superior quality over CoSQA. Models trained on CoSQA+ exhibit improved performance. We provide the code and data at https://github.com/DeepSoftwareAnalytics/CoSQA_Plus.
Index Terms:
software engineering, information search and retrieval, human-computer interaction.
11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
I Introduction
@@
Semantic code search is to retrieve codes that match a given natural language query, which is important for accelerating software development [1, 2, 3, 4, 5]. Programmers can describe their needs in natural language, such as implementing specific algorithms, using particular APIs, or solving complex problems [6]. Given the query, semantic code search methods then retrieve relevant code snippets from the Internet or code repositories. Existing code search benchmarks are designed for one-to-one query-code matching tasks and are typically evaluated using Mean Reciprocal Rank (MRR) [7, 8, 9, 10, 11, 12]. According to our survey of 102 experienced Python programmers with an average of over 1 year of experience, they perform code searches approximately 8.31 times per programming day on average. For each query, they typically reference 2.64 code examples to find the necessary information. These findings highlight a key need in creating a multi-choice code search benchmark: effectively pairing real-world queries with multiple correct and relevant code snippets.
However, existing approaches face significant challenges in scalability, annotation quality, and evaluation metrics:
  1. C1
    Scalability Issues in Human-Based Benchmarking. Conventional benchmarks like CoSQA [8] and CodeSearchNet (CSN) [7] rely on expert human annotators to evaluate query-code pairs. While this approach captures human judgment, manual annotation requires significant time and resources, limiting dataset scalability. The labor-intensive nature of this process makes it impractical for creating large-scale benchmarks needed for robust code search evaluation.
  2. C2
    Accuracy Limitations in Annotation Methods. Both human and LLM-based annotation approaches suffer from accuracy issues. Human annotators typically assess code matching based on semantic understanding rather than functional verification through test execution, which may lead to inaccurate annotations. While Li et al. leverage LLMs to address scalability issues [9], these models struggle to reliably determine whether code functionally satisfies query requirements, as they tend to focus on semantic similarity rather than execution logic. This fundamental limitation affects the quality of both human and LLM-based annotations [13].
  3. C3
    Inadequate Evaluation Metrics for Code Search. Most code search benchmarks employ metrics like MRR or Pass@k [10] which primarily evaluate single-result retrieval performance. This overlooks the reality that queries often have multiple valid matching codes, and users frequently need multiple reference implementations. While alternative metrics like MAP and Recall@k [7] exist for evaluating multiple code matches, they underemphasize the importance of top-ranked results. This is particularly problematic as our survey reveals users heavily rely on first-page search results, highlighting the need for better evaluation metrics.
@@
To address these limitations, we present a new benchmark CoSQA+, with 412,080 pairs of web queries from CoSQA and repository code curated from CodeSearchNet, specifically designed for multi-choice code search. To construct high-quality pairs and achieve a high coverage rate of positive pairs, we employ multiple models to select the top 20 codes with the highest similarity for each query, followed by our novel test-driven agents - a fully automated pipeline comprising a preliminary screener, a test program generator, a test executor, a bug fixer, and a final arbiter. The label indicates whether each code snippet exactly matches its corresponding query requirements. Furthermore, we introduce the metric Normalized Discounted Cumulative Gain at rank 10 (NDCG@10) to code search tasks [14]. We use it to measure 8 code search methods, 7 of which are embedding-based.
@@
To assess the efficacy of our proposed annotation approach, we establish a rigorous evaluation framework using 400 randomly selected query-code pairs from CoSQA+. Three Python experts first design test programs for these pairs to establish ground truth annotations. We then compare the accuracies among test-driven agents, python expert annotators without tests, and a single LLM. Results demonstrate that the test-driven agents achieve superior performance with an accuracy of 92.0%. Notably, 83.67% of the test programs automatically generated by test-driven agents are executable. Furthermore, using identical data sources (queries from CoSQA and code from CodeSearchNet), we evaluate CoSQA+ against CoSQA as fine-tuning corpora for three models: CodeBERT [^1] The version of CodeBERT used in this paper is fine-tuned on the code search task with CodeSearchNet, not the original version., UniXcoder, and CodeT5+ embeddings. Across all three models, fine-tuning on CoSQA+ consistently yields superior performance on the CSN99 Python benchmark, as measured by both NDCG@10 and MRR metrics.
@@
  • To the best of our knowledge, we develop the first Multiagent annotator with 92.0% accuracy for code search dataset.
@@
II Motivation Survey
Do developers really need multi-choice code search? To answer this question, we conduct a survey to investigate the code search behaviors and preferences of Python developers, with a particular focus on understanding how developers interact with multiple code examples during their search process.Fig. 1 provides a multi-choice code search example
Figure 1: An example of multi-choice code search. The query is vague in its definition of "illegal character," allowing eight candidate code snippets to match under different interpretations.
Methodology
The survey targets 102 participants, predominantly undergraduate students (89.22%) with a smaller proportion of master’s students (8.82%). Most of the participants (92.16%) have formal computer science education, and their experience in Python programming varies. The majority (66.67%) have between 1 and 5 years of programming experience, while 30.39% have less than one year of experience. The participants are selected from programmers using Python as their primary programming language (52.94%). The survey is conducted anonymously through an online platform. No personally identifiable information is collected, and participants are informed that their responses will be used for research purposes only.
Results and Analysis
Our survey reveals several significant patterns in code search behavior. Most notably, when 66.67% typically needing 2-3 code examples to understand and implement a solution. This suggests that the current code search task, which typically prioritizes a single result, may need to evolve to provide multiple relevant code snippets. Participants evaluate search results based on several key factors, primarily focusing on implementation relevance (65.69%) and solution differences (56.86%). These preferences emphasize that developers do not just want the closest match to their query, but also value diverse approaches that help them understand the problem and solution better. When presented with multiple code examples, they particularly value the clarity of implementation differences and the completeness of edge case handling.
III Related Work
III-A Code Search Datasets
Besides our work, there are other text-to-code datasets and benchmarks in the realm of code search. They can be categorized into three types based on the origin of the queries: doc-code datasets, question-code datasets, and query-code datasets.
CodeSearchNet [7], CoDesc [15], PyMT5 [16], and XLCoST [17] are doc-code datasets, which contain massive code snippets from public repositories with corresponding comments. They lay foundational training grounds for code search models to grasp basic code-comment relationships. CodeSearchNet [7] dataset comprises a large-scale collection of query-code pairs extracted from GitHub repositories. CoDesc [15] is a comprehensive dataset featuring 4.2 million Java methods paired with natural language descriptions, meticulously filtered to minimize noise. PyMT5, on the other hand, offers a vast collection of 26 million Python methods and 7.7 million method-docstring pairs, enabling bidirectional translation between code and docstrings. XLCoST [17] stands out as the largest multilingual parallel dataset, encompassing 8 languages. It supports 10 cross-lingual code tasks, such as translation and search, and provides essential baselines for model development in diverse programming contexts.
StaQC [11], CoNaLa [18], SO-DS [19], CS1QA [20], and ProCQA [21] are question-code datasets, which contain code answers with corresponding questions from Stack Overflow. They challenge models to excel in practical problem-solving scenarios by understanding complex developer inquiries. StaQC [11] is a large dataset automatically mined from Stack Overflow, which includes multiple languages but varies in code quality. CoNaLa [18] is a curated dataset derived from Stack Overflow, comprising 2,379 training and 500 test examples of natural language-code pairs, alongside a large automatically-mined dataset of 600k examples, designed for tasks like code synthesis, retrieval, and summarization. SO-DS [19], a high-quality snippet collection mined from highly upvoted Stack Overflow posts, focuses on annotated code search, offering 12,137 snippets and 7,674 unique descriptions, with 2,225 annotated queries for validation and testing, specifically tailored for improving code retrieval systems by leveraging natural language descriptions. CS1QA [20], a dataset for code-based question answering in programming education, comprises 9,237 annotated question-answer pairs from Python chat logs, alongside 17,698 unannotated examples, designed to benchmark models on code comprehension and question answering in an educational context. ProCQA [21] introduces a large-scale community-based programming question answering dataset, which is mined from StackOverflow with strict filtering strategies for quality and fairness control. This benchmark adopts the code-mixing dataset, which is more aligned with real-world scenarios. However, the data source is still limited, and the criteria for judging whether the answer matches the question is still vague.
CoSQA [8], CSN99 of CodeSearchNet [7], WebQueryTest of CodeXGLUE [10], and Query4Code [9] are query-code datasets, which contain real human queries from web and code snippets matched with them. They are crucial for training and testing models’ capabilities in handling authentic, diverse search queries. CoSQA [8] introduces a dataset of 20,604 human-annotated natural language query-code pairs, enhancing semantic matching for code search, and proposes CoCLR, a contrastive learning method that boosts code question answering accuracy by 10.5% when combined with CodeBERT. CSN99, part of the CodeSearchNet Challenge [7], offers 99 natural language queries with approximately 4k expert-annotated relevance judgments, facilitating robust evaluation of code search models. The queries of CoSQA and CSN99 are collected from Microsoft Bing search engine. WebQueryTest [10] addresses the gap between real user search queries and existing code search datasets by providing a testing set of about 1k real web queries paired with Python code functions, classified for code-search intent and query-code relevance, with annotations from 13 proficient developers. Query4Code leverages Large Language Models (LLMs) to annotate datasets, primarily relying on the semantic information embedded within the code [9]. However, it lacks test cases or validation mechanisms to ensure the accuracy and reliability of the annotations.
Based on these datasets, COIR [22] emerges as a robust and comprehensive benchmark specifically designed for code retrieval. COIR comprises ten curated datasets spanning eight retrieval tasks across seven diverse domains, offering a versatile evaluation platform for state-of-the-art retrieval models. Its user-friendly Python framework, compatible with popular benchmarks like MTEB [23] and BEIR [24], facilitates seamless integration into research workflows.
Despite the advancements of these code search datasets, several limitations persist. Doc-code datasets oversimplify the search task by focusing on direct comment-to-code mappings, lacking the complexity of real-world queries. Question-code datasets are not designed for code search tasks, emphasizing question-answer context over direct code retrieval. Query-code datasets, though highly relevant, are constrained by their difficulty in constructing and annotating at scale.
III-B Code Search Methods
The existing methods for code search can mainly be divided into two mainstreams: information retrieval (IR) based methods and deep learning (DL) based models. IR-based methods extract keywords from queries and search for these keywords in the code base to retrieve the most relevant code snippets [25, 26, 27]. However, standard information retrieval methods have drawbacks in the code search domain, as they treat codes as text while programming languages and natural languages differ greatly. Fortunately, DL-based methods receive great attention in NLP and achieve significant breakthroughs in code search research. The CodeSearchNet [7] dataset paves the way for CodeBERT [12] and GraphCodeBERT [28], both of which are bimodal pre-trained BERT [29] based models that leverage Masked Language Modeling (MLM) and Replaced Token Detection (RTD) [30] objective. UniXcoder [31] unifies three pre-training designs into one architecture and utilizes AST structure and code comment to enhance the cross-modal alignment. Recent T5 [32] based models like CodeT5 [33] and CodeT5+ [34] show outstanding performance in code search, utilizing the text-to-text transformer architecture and pre-training on large code datasets. While the two models achieve high performances in the CSN benchmark, GraphCodeBERT [28] demonstrates superior results by harnessing data flow encoded in the Abstract Syntax Tree (AST) of code to enrich the structural information during pre-training. Recent advancements in text embedding models have significantly enhanced the performance of code search methods, particularly in multilingual and long-context retrieval tasks. Among these, jina-embeddings-v3 [35] stands out as a state-of-the-art model with 570 million parameters, achieving superior performance on the MTEB benchmark. Its innovative use of task-specific Low-Rank Adaptation (LoRA) [36] adapters enables high-quality embeddings for various tasks, including query-document retrieval and text matching, while supporting context lengths of up to 8192 tokens. This model outperforms both proprietary embeddings like OpenAI and Cohere. multilingual-e5-large [37, 38], which is trained on 1 billion multilingual text pairs and fine-tuned on labeled datasets, is efficient, especially in reducing embedding dimensions without compromising quality, thanks to Matryoshka Representation Learning [39]. But it does not match the flexibility and performance of jina-embeddings-v3. On the other hand, all-MiniLM-L12-v2 [40] and all-mpnet-base-v2 [41] represent efficient, compressed models derived from larger pre-trained Transformers. The all-MiniLM-L12-v2 model, based on MiniLM [42], employs deep self-attention distillation to retain high accuracy with reduced parameters, making it suitable for latency-sensitive applications. Similarly, all-mpnet-base-v2, built on MPNet [43], combines the strengths of BERT and XLNet [44] by leveraging masked and permuted language modeling, achieving competitive results on downstream tasks [45]. While these models are effective for general-purpose embeddings, they lack the multilingual and long-context capabilities of jina-embeddings-v3 and multilingual-e5-large, limiting their applicability in complex code search scenarios.
III-C Test-Driven Agents
@@
Traditional test generation approaches rely heavily on human expertise, but recent work demonstrates the potential of automated test agents. LLM-based agents have shown promising results in various testing scenarios. For instance, in the domain of deep learning libraries, agents capable of generating diverse and meaningful test inputs are proposed [46, 47], alongside other studies exploring LLM-based agents for specialized tasks. For test oracle generation, few-shot prompting is used to generate context-aware assertions for test oracle generation [48, 49], achieving a high accuracy. In program repair, some studies [50, 51, 52, 53, 54] demonstrate the effectiveness of integrating LLMs with iterative validation workflows [55]. However, existing test-driven agents primarily focus on validating specific code functionality without addressing the semantic challenges in code search systems, where our test-driven annotation approach bridges this gap through systematic validation of query-code relevance [56].
Figure 2: The construction process of CoSQA+. (a) Queries sourced from CoSQA and code snippets filtered from CodeSearchNet. (b) Candidate pairs selected based on query-code cosine similarity, computed using a multi-model approach. (c) Pipeline for annotation by test-driven agents.
IV CoSQA+
@@
Fig. 2 illustrates the benchmark construction process for CoSQA+. The process starts with query and code collection, in which we leverage the queries from CoSQA and collect python code snippets filtered from CodeSearchNet. In candidate pairs construction, collected code snippets and queries are embedded as vectors to calculate similarity. For each query, the top 20 query-code pairs with the highest similarity are selected as candidates for potential matches. In model annotation, each query-code pair is annotated by test-driven agents with “1” or “0”, indicating whether the code snippet exactly matches the query. These labeled pairs are then added to CoSQA+.
IV-A Query and Code Collection
@@
This section details the reasons why we select CoSQA as the query source and the CodeSearchNet Corpus [7] serves as the code sources.
@@
Collected from Microsoft’s Bing search engine logs, CoSQA queries are based on real user queries for python codes [8]. We use the queries as the source for the following reasons: (1) These queries are drawn directly from real-world programming questions, making them representative of actual developer needs and search patterns. For the practical usability of neural code search, such authentic queries are crucial to creating high-quality datasets [57]. (2) Through rigorous filtering and validation processes, the queries demonstrate strong consistency in format and clear programming intent. This careful curation helps maintain high data quality and reduces noise in the training process [58]. By training on these curated real-world queries, code search models can better align with how developers naturally express their programming needs.
@@
To the best of our knowledge, CodeSearchNet is the only origin dataset specifically designed for Python functions in the domain of code search. While other datasets, such as CoSQA, also feature Python codes, their code bases are derived from CodeSearchNet. The CodeSearchNet Corpus contains about 6.4 million functions from open-source code spanning six programming languages [7]. According to our survey, code searchers prefer function codes with clear input and output. These codes are high quality with higher readability, testability and maintainability [59]. To extract these high-quality codes from CodeSearchNet’s 1,156,085 Python functions, we filter out 653,994 functions using a rule-based approach. This involves parsing the code into an Abstract Syntax Tree (AST), identifying function nodes with at least one input parameter and a valid return statement. Such testable functions lay the foundation for test-driven agents to design test program and improve dataset quality.
IV-B Candidate Pairs Construction